OpenStudy (anonymous):

Calculate dy/dx. You need not expand your answer. y = 3x^2 − 9x + 13/2x + 4

OpenStudy (amistre64):

\[[\frac{t}{b}]'=\frac{bt'-b't}{b^2}\]

OpenStudy (anonymous):

\[y'=6 x-\frac{5}{2} \]